after taking 3 math quizzes elisa has an average of 89. what must she score on the fourth quiz t raise her average to 91.
Answer by stanbon(48558)

after taking 3 math quizzes elisa has an average of 89. what must she score on the fourth quiz t raise her average to 91.
-------------
If she averaged 89 on 3 quizzes she has 3*89 = 267 points.
Let x be the score she needs to average 91:
(267+x)/4 = 91
267+x = 364
x = 97 (score she needs on the 4th quize)
============
Cheers,
Stan H.